OpenStudy (anonymous):

Using the balanced equation 2Na(s) + Cl2(g) --> 2NaCl(s), what is the volume of Cl2 gas needed to form 23 grams of NaCl at a temperature of 37C and 689 torr?

OpenStudy (aaronq):

first, find how many moles 23 g of NaCl are.

OpenStudy (aaronq):

After you wanna use the coefficients to relate the 2 subtances. Using the stoichiometric coefficients to find moles produced. -------------------------------------------------------- set it up like this: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) from here you can solve for the moles of anyone substance. -------------------------------------------------------- For example: if you have moles of B= 2, how many moles of C did you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_c=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) -------------------------------------------------------- Then use the ideal gas law: PV=nRT

OpenStudy (anonymous):

Thank you this is a big help. Getting started is usually my problem.

OpenStudy (aaronq):

no problem !